OpenStudy (rachie19):

3. A weather balloon is released at sea level (P=1.00 atm, T = 298 K, V = 25 L). Assuming no loss of gas as the balloon rises, determine the pressure in the upper atmosphere when the temperature is 255K and the volume has increased by 15% over what the volume was at sea level.

OpenStudy (h2so4):

You use (P1*V1)/T1 = (P2*V2)/T2, no? Put the numbers in P1, V1, T1, V2, T2 to find P2. https://en.wikipedia.org/wiki/Combined_gas_law

OpenStudy (h2so4):

Oh right. Convert the units of yours into the ones shown on Wikipedia.

OpenStudy (rachie19):

so (1.00atm)(25l)/298k=p2(18.75)/255